# Efficient Compositional Multi-tasking for On-device Large Language Models

### v1

@@ Abstract
Adapter parameters provide a mechanism to modify the behavior of machine learning models and have gained significant popularity in the context of large language models (LLMs) and generative AI. These parameters can be merged to support multiple tasks via a process known as task merging. However, prior work on merging in LLMs, particularly in natural language processing, has been limited to scenarios where each test example addresses only a single task. In this paper, we focus on on-device settings and study the problem of text-based compositional multi-tasking, where each test example involves the simultaneous execution of multiple tasks. For instance, generating a translated summary of a long text requires solving both translation and summarization tasks concurrently. To facilitate research in this setting, we propose a benchmark comprising four practically relevant compositional tasks. We also present an efficient method (Learnable Calibration) tailored for on-device applications, where computational resources are limited, emphasizing the need for solutions that are both resource-efficient and high-performing. Our contributions lay the groundwork for advancing the capabilities of LLMs in real-world multi-tasking scenarios, expanding their applicability to complex, resource-constrained use cases.
@@ 1 Introduction
Large language models (LLMs) have revolutionized natural language processing (NLP) by demonstrating remarkable capabilities across a wide range of text-based tasks Zhao et al. (2023); Minaee et al. (2024). As foundational models, LLMs can be fine-tuned to excel in various downstream applications, such as question answering Sticha et al. (2024); Zhang et al. (2024a), summarization Liu et al. (2024b); Zhang et al. (2024b), translation Zhu et al. (2023); He et al. (2024), and beyond Shu et al. (2024); Severyn et al. (2021). A common approach to fine-tuning or adapting LLMs is through parameter-efficient fine-tuning (PEFT), where relatively few additional parameters, often referred to as adapters, are inserted into the model Han et al. (2024); Ding et al. (2022). Among these methods, Low-Rank Adapters (LoRA) Hu et al. (2022) have gained prominence due to their efficiency and ease of integration. LoRA trains low-rank matrices inserted into the LLM’s layers, enabling efficient adaptation to specific tasks. Typically, models are fine-tuned for one task at a time, producing task-specific models or experts that can be merged to obtain a final model supporting multiple tasks Wortsman et al. (2022); Ilharco et al. (2023). A task may correspond to, for example, (i) different domains of expertise (e.g., math versus biology Q&A), (ii) different downstream tasks (e.g., summarization versus translation), or (iii) different optimization convergence points (e.g., results from separate training runs). We focus on LLM customization for different downstream tasks.
@@ 2 Related Work
On-device LLMs. LLMs often contain billions of parameters, necessitating significant resources, such as multiple high-end GPUs, for inference Borzunov et al. (2024). However, many valuable LLM use cases involve sensitive data stored on resource-constrained devices, making it desirable to perform computations locally and avoid transferring data to remote servers Dhar et al. (2021). For example, users may wish to summarize private conversations or generate personalized replies while maintaining data privacy. To support such use cases, smaller LLMs (e.g., 1–3 billion parameters) have been developed for on-device deployment. These models leverage compression techniques and smaller parameter sizes to enable efficient on-device inference. Prominent examples include LLaMA 3.2 1B Dubey et al. (2024), Qwen2.5 1.5B Yang et al. (2024a); Qwen Team (2024), and StableLM2 1.6B Bellagente et al. (2024). Single-task LoRAs are typically stored on-device to enable an LLM to support the individual tasks, rather than relying on instruction following Gunter et al. (2024). On-device LLMs also have a limited context window size, making in-context learning less suitable Dong et al. (2024).
@@ 3 Benchmark
Summarization + Tone Adjustment. We change the tone of the ground-truth summaries available in DialogSum via the publicly available RedPajama-INCITE-Base 3B model Weber et al. (2024) fine-tuned specifically for tone adjustment and paraphrasing Utsav (2023).
@@
Reply + Tone Adjustment. Tone-adjusted replies are generated from the Synthetic Persona Chat dataset using the RedPajama-INCITE-Base 3B model, i.e., the model specialized for tone adjustment and paraphrasing.
@@
[…]
@@ 4 Method
Existing approaches that can be used for compositional multi-tasking in on-device LLMs are either inefficient or have low performance, as we show in our analysis. How can we achieve both efficiency and good performance so that we can target on-device applications where computational resources and storage are restricted? We propose a new method that addresses this challenge. In particular, we develop an efficient solution that achieves comparable or superior performance to inefficient baselines, such as the multi-step application of adapters or a new adapter trained specifically for each compositional task (referred to as a "joint-expert" adapter).
@@ 4.1.1 Baseline and Merging Strategies
We also consider model merging strategies, which have shown success in standard multi-tasking where test examples involve a single task at a time. In particular, we consider the following merging strategies applied to LoRA adapters: 1) Linear, i.e., Model Soup Wortsman et al. (2022); 2) Concatenation Mangrulkar et al. (2022); 3) TIES Yadav et al. (2024); 4) DARE Yu et al. (2024); 5) Slerp White (2016); 6) LoraHub Huang et al. (2024); 7) LM-Cocktail Xiao et al. (2024); 8) DAM Gauthier-Caron et al. (2024) and ZipLoRA Shah et al. (2024) adapted to our use cases. These methods are typically designed for merging entire models rather than LoRA adapters. We adapt them to merge LoRAs and evaluate their performance on compositional tasks.
@@ 5.2 Main Results
Table 2: Benchmark of compositional multi-tasking. Test results reported as % ($\uparrow$) and averaged across models and languages or tones. Our Learnable Calibration methods achieve comparable performance to inefficient baselines while being significantly more efficient in terms of inferences and storage. Similarly, fast baselines, such as various merging strategies, typically fail in compositional multi-tasking. The bottom block includes results for a version of Learnable Calibration where additional parameters are shared across all four tasks. The best three methods from the main set of results are in bold.
@@ 5.3 Analyses
We have studied a wide range of questions within the additional analyses, for example sharing of parameters across tasks, efficiency, scaling to models of various sizes and compositional multi-tasking with three tasks.
@@
Efficiency Analysis. We compare the efficiency of our solutions against the two well-performing but inefficient baselines in Table 5.3. Our solutions require only a minimal number of additional parameters, amounting to approximately 0.08–0.56% of the parameters of a joint-expert LoRA. The resulting additional storage on disk is less than 0.5 MB, making our solutions suitable for on-device deployment.
Table 3: Efficiency of well-performing approaches. Our methods require only 0.08–0.56% of additional parameters/storage, depending on the variation (averaged across models). Baselines not reported here, such as Main-task LoRA and Linear Merge, are efficient (i.e., single inference pass, zero additional parameters/storage) but exhibit significantly lower performance.
Method # of
Additional
Additional
Storage
Multi-step LoRAs 2×2$\times$2 × 0 00.00MB
Joint-expert LoRA 1×1$\times$1 × 030M 57.10MB
[…]
Learnable Calibration++ 1×1$\times$1 × 166K 00.32MB
@@
Scaling Analysis. LLMs designed for mobile device deployment are available in various sizes, including newer models with up to 3B parameters Gunter et al. (2024); Carreira et al. (2023). We analyze the scalability of our solution across different model sizes in Figure 6. For this analysis, we use the Qwen2.5 model, which is available in three sizes suitable for on-device deployment: 0.5B, 1.5B, and 3B parameters. The results demonstrate that our solutions consistently perform well across all model sizes. Additionally, as expected, larger models typically achieve stronger performance.
Figure 6: Evaluation across model sizes. Evaluation of compositional multi-tasking on the cross-lingual summarization task for the Qwen2.5 model across different sizes: 0.5B, 1.5B, and 3B. Test results for ROUGE-L and LLM Judge score (%, $\uparrow$) are averaged across languages. Our solutions consistently demonstrate strong performance across all model sizes, with performance improving as model size increases.
Out-of-Domain Generalization. We evaluate the generalization of LoRAs and additional parameters trained on the DialogSum dataset by testing them on the SAMSum dataset Gliwa et al. (2019). Our aim is to assess whether our method maintains strong performance and improvements when applied to data from a different domain. Specifically, DialogSum features spoken-language conversations from daily life scenarios, while SAMSum includes written-style dialogues from online interactions. For instance, DialogSum contains sentences like “#Person_1#: Good morning. I wonder whether you have got an answer from your superior.,” whereas SAMSum includes examples such as “Leo: BTW what are those pics?
Table 5 presents results for both in-domain and out-of-domain settings using the cross-lingual English-to-Spanish summarization task. The evaluation spans all three models considered earlier. The results demonstrate that the trained LoRAs and additional parameters deliver strong performance even under mild domain shifts, such as variations in conversation style.
Table 5: Comparison of in-domain and out-of-domain settings. Test scores reported as % ($\uparrow$) and averaged across models for the cross-lingual English-to-Spanish summarization task. The approaches work well even in the presence of mild domain shift.
In-domain Out-of-domain
R-L LLM-J R-L LLM-J
Zero-shot 09.59 00.20 10.58 03.54
Main-task LoRA 14.25 03.09 15.97 09.89
Auxiliary-task LoRA 14.42 00.16 19.67 01.14
Linear merge 15.60 00.18 18.30 01.26
[…]
Joint-expert LoRA 23.47 52.51 25.38 56.33
[…]
Learnable Calibration++ 30.98 67.98 28.97 59.91
Compositional Multi-tasking with Three Tasks. We demonstrate that combinations involving three tasks can also be effectively addressed using our Learnable Calibration framework. Specifically, we evaluate the joint tasks of summarization, professional tone adjustment, and translation from English to Spanish, German, or French. We created data for this task by converting ground-truth summaries into a professional tone and then translating them into multiple languages. The same models for tone adjustment and translation were applied in this setup. Given the large number of possible combinations, we focus on one tone and main task for simplicity. Additionally, we evaluate only a subset of the most relevant approaches for this scenario.
Table 6: Compositional multi-tasking with three-way tasks. Combination of summarization, professional tone adjustment, and translation to various languages. Test scores reported as % ($\uparrow$) and averaged across languages and models. Our solution also works well for three-way compositional tasks.
R-L LLM-J
Zero-shot 07.35 00.03
Summarization LoRA 09.41 00.00
Translation LoRA 10.95 00.00
Tone adj. LoRA 08.39 00.01
Linear merge 09.89 00.01
[…]
Joint-expert LoRA 14.35 12.75
[…]
Learnable Calibration++ 21.42 39.16
We report the results of our compositional multi-tasking analysis with three-way tasks in Table 6. The findings show: 1) inefficient baselines outperform simple baselines in this more challenging scenario, 2) our Learnable Calibration approach outperforms the inefficient baselines both in terms of accuracy and footprints, and 3) Learnable Calibration++ performs better than its simpler variant. While the overall performance is lower than that for the simpler combination of summarization and translation, these results confirm our framework can successfully handle compositions involving more than two tasks, further proving its flexibility.
Interpretation of the Impact of Different Solutions. ROUGE scores measure the overlap between generated and reference texts but can be challenging to interpret. To better understand these differences, we conducted a qualitative analysis, detailed in Appendix D, and summarized here.
6 Discussion
While LLMs typically incorporate built-in safeguarding mechanisms, these safeguards may be weakened when models are merged Hammoud et al. (2024) or adapted for compositional multi-tasking. It is crucial to ensure safeguards remain effective in these scenarios before deployment.
Future research could explore the interplay between compositional multi-tasking and safeguarding mechanisms, with a focus on developing robust, high-performing safeguards for the proposed setup. Addressing these challenges will be essential for the responsible advancement and deployment of compositional multi-tasking capabilities.
@@ Appendix C Extended Evaluation with Standard Metrics
Appendix D Qualitative Analysis
In this section, we perform qualitative analysis using test examples with ROUGE scores close to the average for each method. We compare the performance of three approaches—zero-shot, main-task LoRA, and Learnable Calibration++—to illustrate varying levels of success. For this analysis, we use the Qwen2.5 1.5B model with either Spanish as the target language or professional tone as the target tone.
Figures A4, A5 and A6 highlight common issues encountered by zero-shot and main-task LoRA methods:

### v2

@@ Abstract
Adapter parameters provide a mechanism to modify the behavior of machine learning models and have gained significant popularity in the context of large language models (LLMs) and generative AI. These parameters can be merged to support multiple tasks via a process known as task merging. However, prior work on merging in LLMs, particularly in natural language processing, has been limited to scenarios where each test example addresses only a single task. In this paper, we focus on on-device settings and study the problem of text-based compositional multi-tasking, where each test example involves the simultaneous execution of multiple tasks. For instance, generating a translated summary of a long text requires solving both translation and summarization tasks concurrently. To facilitate research in this setting, we propose a benchmark comprising four practically relevant compositional tasks. We also present an efficient method (Learnable Calibration) tailored for on-device applications, where computational resources are limited, emphasizing the need for solutions that are both resource-efficient and high-performing. Our contributions lay the groundwork for advancing the capabilities of LLMs in real-world multi-tasking scenarios, expanding their applicability to complex, resource-constrained use cases. Project page: https://ondrejbohdal.github.io/CompositionalMultitaskingLLMs.
@@ 1 Introduction
Large language models (LLMs) have revolutionized natural language processing (NLP) by demonstrating remarkable capabilities across a wide range of text-based tasks Zhao et al. (2023); Minaee et al. (2024). As foundational models, LLMs can be fine-tuned to excel in various downstream applications, such as question answering Sticha et al. (2024), summarization Liu et al. (2024b), translation Zhu et al. (2023), and beyond Shu et al. (2024); Severyn et al. (2021). A common approach to fine-tuning or adapting LLMs is through parameter-efficient fine-tuning (PEFT), where relatively few additional parameters, often referred to as adapters, are inserted into the model Han et al. (2024); Ding et al. (2023). Among these methods, Low-Rank Adapters (LoRA) Hu et al. (2022) have gained prominence due to their efficiency and ease of integration. LoRA trains low-rank matrices inserted into the LLM’s layers, enabling efficient adaptation to specific tasks. Typically, models are fine-tuned for one task at a time, producing task-specific models or experts that can be merged to obtain a final model supporting multiple tasks Wortsman et al. (2022); Ilharco et al. (2023). A task may correspond to, for example, (i) different domains of expertise (e.g., math versus biology Q&A), (ii) different downstream tasks (e.g., summarization versus translation), or (iii) different optimization convergence points (e.g., results from separate training runs). We focus on LLM customization for different downstream tasks.
@@ 2 Related Work
On-device LLMs. LLMs often contain billions of parameters, necessitating significant resources, such as multiple high-end GPUs, for inference Borzunov et al. (2024). However, many valuable LLM use cases involve sensitive data stored on resource-constrained devices, making it desirable to perform computations locally and avoid transferring data to remote servers Dhar et al. (2021). For example, users may wish to summarize private conversations or generate personalized replies while maintaining data privacy. To support such use cases, smaller LLMs have been developed for on-device deployment. These models leverage compression techniques and smaller parameter sizes to enable efficient on-device inference. Prominent examples include LLaMA 3.2 1B Dubey et al. (2024), Qwen2.5 1.5B Yang et al. (2024a); Qwen Team (2024), and StableLM2 1.6B Bellagente et al. (2024). Our work assumes that model size is the key constraint for on-device deployment, as current mobile-compatible LLMs typically range from 1–3B parameters. Single-task LoRAs are typically stored on-device to enable an LLM to support the individual tasks, rather than relying on instruction following Gunter et al. (2024). On-device LLMs also have a limited context window size, making in-context learning less suitable Dong et al. (2024).
@@ 3 Benchmark
Summarization + Tone Adjustment. We change the tone of the ground-truth summaries available in DialogSum via the publicly available RedPajama-INCITE-Base 3B model Weber et al. (2024) fine-tuned specifically for tone adjustment and paraphrasing Utsav (2023). We use the corresponding prompt that was predefined for tone adjustment.
@@
Reply + Tone Adjustment. Tone-adjusted replies are generated from the Synthetic Persona Chat dataset using the RedPajama-INCITE-Base 3B model, i.e., the model specialized for tone adjustment (adj.) and paraphrasing.
@@
$\text{W-R}=\frac{\text{ROUGE-1}}{6}+\frac{\text{ROUGE-2}}{3}+\frac{\text{ROUGE-3}}{2}.$ (2)
@@ 4 Method
Existing approaches that can be used for compositional multi-tasking in on-device LLMs are either inefficient or have low performance, as we show in our analysis. How can we achieve both efficiency and good performance so that we can target on-device applications where computational resources and storage are restricted? We propose a new method that addresses this challenge. In particular, we develop an efficient solution that achieves comparable or superior performance to inefficient baselines, such as the multi-step application of adapters or a new adapter trained specifically for each compositional task (referred to as a “joint-expert” adapter).
@@ 4.1.1 Baseline and Merging Strategies
We also consider model merging strategies, which have shown success in standard multi-tasking where test examples involve a single task at a time. In particular, we consider the following merging strategies applied to LoRA adapters: 1) Linear, i.e., Model Soup Wortsman et al. (2022); 2) Concatenation Mangrulkar et al. (2022); 3) TIES Yadav et al. (2024); 4) DARE Yu et al. (2024); 5) Slerp White (2016); 6) LoraHub Huang et al. (2024); 7) LM-Cocktail Xiao et al. (2024); 8) DAM Gauthier-Caron et al. (2024) and ZipLoRA Shah et al. (2024) adapted to our use cases. These methods are typically designed for merging entire models rather than LoRA adapters. We adapt them to merge LoRAs and evaluate their performance on compositional tasks. Note that joint-expert LoRA, LoraHub, LM-Cocktail and DAM/ZipLoRA approaches leverage extra compositional task data.
@@ 4.3 Variations of Learnable Calibration
The on-device pipeline of our system involves two steps: (i) Computing the merged, calibrated LoRA $\Delta W^{c}$ via matrix operations (6) or (7), which is extremely fast and negligible in runtime; (ii) Loading $\Delta W^{c}$ onto the model, which results in inference latency and throughput identical to standard LoRAs. Our method is compatible with existing frameworks such as Android AI Core Burke (2023) and Apple Intelligence Gunter et al. (2024), making it readily deployable in real-world systems.
@@ 5.2 Main Results
Table 2: Benchmark of compositional multi-tasking. Test results reported as % ($\uparrow$) and averaged across models and languages or tones. Our Learnable Calibration methods achieve comparable performance to inefficient baselines while being significantly more efficient in terms of inferences and storage. Similarly, fast baselines, such as various merging strategies, typically fail in compositional multi-tasking. The metrics include ROUGE-L (R-L), Weighted ROUGE (W-R) and LLM judge (LLM-J) scores. The Efficient? column captures both runtime and storage efficiency. The bottom block includes results for a version of Learnable Calibration where additional parameters are shared across all four tasks. The best three methods from the main set of results are in bold.
@@ 5.3 Analyses
We have performed various further analyses, including ones about sharing of parameters across tasks, efficiency, benefit of using existing adapters and analysis of changes in the update matrices.
@@
Efficiency Analysis. We compare the efficiency of our solutions against the two well-performing but inefficient baselines in Table 3. Our solutions require only a minimal number of additional parameters, amounting to approximately 0.08–0.56% of the parameters of a joint-expert LoRA. The resulting additional storage on disk is less than 0.5 MB, making our solutions suitable for on-device deployment.
Table 3: Efficiency of well-performing approaches. Our methods require only 0.08–0.56% of additional parameters/storage, depending on the variation (averaged across models). Baselines not reported here, such as Main-task LoRA and Linear Merge, are efficient (i.e., single inference pass, zero additional parameters/storage) but have significantly lower performance.
Method # of Inferences Additional Parameters Additional Storage
Multi-step LoRAs $2\times$ 0 00.00MB
Joint-expert LoRA $1\times$ 030M 57.10MB
Learnable Calibration $1\times$ 023K 00.05MB
Learnable Calibration++ $1\times$ 166K 00.32MB
@@
Further Analyses. Notably, we show Learnable Calibration 1) works well for models of various sizes such as from 0.5B to 3B (Appendix D), 2) is able to handle domain shift (Appendix E), and 3) obtains strong performance also for three-way compositional tasks (Appendix F). We also perform qualitative analysis in Appendix G. The analysis shows there are essentially two behavioral groups. The first group typically does not perform one of the tasks correctly and includes approaches such as zero-shot and existing merging strategies. The second group succeeds in performing both tasks and includes the two inefficient baselines as well as our Learnable Calibration solutions.
@@ 6 Conclusion
Limitations
Ethical Considerations
While LLMs typically incorporate built-in safeguarding mechanisms, these safeguards may be weakened when models are merged Hammoud et al. (2024) or adapted for compositional multi-tasking. It is crucial to ensure that safeguards remain effective in these scenarios before deploying solutions.
Potential future research could explore the interplay between compositional multi-tasking and safeguarding mechanisms, with a focus on developing robust, high-performing safeguards for the proposed setup. Addressing these challenges will be essential for the responsible advancement and deployment of compositional multi-tasking capabilities.
@@ Appendix C Extended Evaluation with Standard Metrics
Appendix D Scaling Analysis
LLMs designed for mobile device deployment are available in various sizes, including newer models with up to 3B parameters Gunter et al. (2024); Carreira et al. (2023). We analyze the scalability of our solution across different model sizes in Figure A4. For this analysis, we use the Qwen2.5 model, which is available in three sizes suitable for on-device deployment: 0.5B, 1.5B, and 3B parameters. The results demonstrate that our solutions consistently perform well across all model sizes. Additionally, as expected, larger models typically achieve stronger performance.
Figure A4: Evaluation across model sizes. Evaluation of compositional multi-tasking on the cross-lingual summarization task for the Qwen2.5 model across different sizes: 0.5B, 1.5B, and 3B. Test results for ROUGE-L and LLM Judge score (%, $\uparrow$) are averaged across languages. Our solutions consistently demonstrate strong performance across all model sizes, with performance improving as model size increases.
Appendix E Out-of-Domain Generalization
We evaluate the generalization of LoRAs and additional parameters trained on the DialogSum dataset by testing them on the SAMSum dataset Gliwa et al. (2019). Our aim is to assess whether our method maintains strong performance and improvements when applied to data from a different domain. Specifically, DialogSum features spoken-language conversations from daily life scenarios, while SAMSum includes written-style dialogues from online interactions. For instance, DialogSum contains sentences like “#Person_1#: Good morning. I wonder whether you have got an answer from your superior.,” whereas SAMSum includes examples such as “Leo: BTW what are those pics?
Table A3 presents results for both in-domain and out-of-domain settings using the cross-lingual English-to-Spanish summarization task. The evaluation spans all three models considered earlier. The results demonstrate that the trained LoRAs and additional parameters deliver strong performance even under mild domain shifts, such as variations in conversation style.
Table A3: Comparison of in-domain and out-of-domain settings. Test scores reported as % ($\uparrow$) and averaged across models for the cross-lingual English-to-Spanish summarization task. The approaches work well even in the presence of mild domain shift.
In-domain Out-of-domain
R-L LLM-J R-L LLM-J
Zero-shot 09.59 00.20 10.58 03.54
Main-task LoRA 14.25 03.09 15.97 09.89
Auxiliary-task LoRA 14.42 00.16 19.67 01.14
Linear merge 15.60 00.18 18.30 01.26
Multi-step LoRA usage 21.71 75.51 26.81 76.35
Joint-expert LoRA 23.47 52.51 25.38 56.33
Learnable Calibration 27.68 61.58 27.78 57.26
Learnable Calibration++ 30.98 67.98 28.97 59.91
Appendix F Compositional Multi-tasking with Three Tasks
We demonstrate that combinations involving three tasks can also be effectively addressed using our Learnable Calibration framework. Specifically, we evaluate the joint tasks of summarization, professional tone adjustment, and translation from English to Spanish, German, or French. We created data for this task by converting ground-truth summaries into a professional tone and then translating them into multiple languages. The same models for tone adjustment and translation were applied in this setup. Given the large number of possible combinations, we focus on one tone and main task for simplicity. Additionally, we evaluate only a subset of the most relevant approaches for this scenario.
Table A4: Compositional multi-tasking with three-way tasks. Combination of summarization, professional tone adjustment, and translation to various languages. Test scores reported as % ($\uparrow$) and averaged across languages and models. Our solution also works well for three-way compositional tasks.
R-L LLM-J
Zero-shot 07.35 00.03
Summarization LoRA 09.41 00.00
Translation LoRA 10.95 00.00
Tone adj. LoRA 08.39 00.01
Linear merge 09.89 00.01
Multi-step LoRA usage 16.50 43.82
Joint-expert LoRA 14.35 12.75
Learnable Calibration 19.19 28.66
Learnable Calibration++ 21.42 39.16
We report the results of our compositional multi-tasking analysis with three-way tasks in Table A4. The findings show: 1) inefficient baselines outperform simple baselines in this more challenging scenario, 2) our Learnable Calibration approach outperforms the inefficient baselines both in terms of accuracy and footprints, and 3) Learnable Calibration++ performs better than its simpler variant. While the overall performance is lower than that for the simpler combination of summarization and translation, these results confirm our framework can successfully handle compositions involving more than two tasks, further proving its flexibility.
Appendix G Qualitative Analysis
ROUGE scores measure the overlap between generated and reference texts but can be challenging to interpret. To better understand these differences, we conducted a qualitative analysis.
More specifically, we perform the qualitative analysis using test examples with ROUGE scores close to the average for each method. We compare the performance of three approaches—zero-shot, main-task LoRA, and Learnable Calibration++—to illustrate varying levels of success. For this analysis, we use the Qwen2.5 1.5B model with either Spanish as the target language or professional tone as the target tone.
Figures A5, A6 and A7 highlight common issues encountered by zero-shot and main-task LoRA methods: